# Restricted $L_{\infty}$-algebras and a derived Milnor-Moore theorem

### v1

@@ 1. Introduction
(Theorem 6.7) Let ${\mathrm{K}}$ be a field. There is a forgetful functor
$\mathrm{Alg}_{{\mathrm{Prim}}{\mathrm{T}}}({\mathrm{Mod}}_{{\mathrm{H}}({\mathrm{K}})}^{\geq 0})\to(\mathsf{s}{\mathrm{Lie}}^{\mathrm{res}}_{\mathrm{K}})_{\infty}$
from the $\infty$-category of restricted Lie algebras in connective ${\mathrm{H}}({\mathrm{K}})$-modules to the $\infty$-category underlying a right induced model structure on the category of simplicial restricted Lie algebras over ${\mathrm{K}}$.
@@ 1.1. Acknowledgements
We thank Lukas Brantner, Gijs Heuts, Ben Knudsen, Markus Spitzweck and Manfred Stelzer for helpful and inspiring discussions about this topic. We are especially grateful to Gijs Heuts for communicating to us the proof of Lemma 5.23.
This work are main parts of my thesis written during my phd 2014-2018 at the university of Osnabrück. I am grateful for this opportunity.
@@ 1.2. Notation and Terminologie
  • $\mathsf{Set}$ for the category of small sets.
  • $\Delta$ for the category of finite, non-empty, partially ordered sets and order preserving maps, whose objects we denote by $[{\mathrm{n}}]=\{0<...<{\mathrm{n}}\}$ for ${\mathrm{n}}\geq 0$.
  • ${\mathrm{{\mathcal{F}}in}}$ for (a skeleton of) the category of finite sets, whose objects we denote by ${\mathrm{n}}=\{1,...,{\mathrm{n}}\}$ for ${\mathrm{n}}\geq 0$.
  • ${\mathrm{{\mathcal{F}}in}}_{*}$ for (a skeleton of) the category of finite pointed sets, whose objects we denote by $\langle{\mathrm{n}}\rangle=\{*,1,...,{\mathrm{n}}\}$ for ${\mathrm{n}}\geq 0$.
@@
  • $\mathsf{Cat}_{\infty}$ for the $\infty$-category of small $\infty$-categories.
  • $\mathsf{Cat}_{\infty}^{{\mathrm{c}}{\mathrm{c}}}$ for the $\infty$-category of large $\infty$-categories with small colimits and small colimits preserving functors.
We often indicate $\infty$-categories of large objects by $\widehat{(-)}$, for example we write $\widehat{{\mathcal{S}}},\widehat{\mathsf{Cat}}_{\infty}$ for the $\infty$-categories of large spaces, $\infty$-categories.
@@
  • $\mathsf{Ho}({\mathcal{C}})$ for its homotopy category,
  • ${\mathcal{C}}^{\triangleleft},{\mathcal{C}}^{\triangleright}$ for the $\infty$-category arising from ${\mathcal{C}}$ by adding an initial, final object, respectively,
@@
We heavily use symmetric $\infty$-operads [23, Definition 2.1.1.10.] and symmetric monoidal $\infty$-categories [23, Definition 2.0.0.7.]. For every $\infty$-operad ${\mathcal{O}}^{\otimes}\to{\mathrm{{\mathcal{F}}in}}_{*}$ we write ${\mathcal{O}}$ for the fiber over $\langle 1\rangle\in{\mathrm{{\mathcal{F}}in}}_{*}$ and call ${\mathcal{O}}$ the underlying $\infty$-category. More generally, we use ${\mathcal{O}}$-operads and ${\mathcal{O}}$-monoidal $\infty$-categories [23, Definition 2.1.2.13.], which for ${\mathcal{O}}^{\otimes}={\mathrm{{\mathcal{F}}in}}_{*}$ give symmetric $\infty$-operads and symmetric monoidal $\infty$-categories, for ${\mathcal{O}}^{\otimes}={\mathrm{Ass}}$ give non-symmetric $\infty$-operads and monoidal $\infty$-categories and for ${\mathcal{O}}^{\otimes}={\mathrm{LM}},{\mathrm{RM}},{\mathrm{BM}}$ give ${\mathrm{LM}},{\mathrm{RM}},{\mathrm{BM}}$-operads and left, right, bitensored $\infty$-categories.
Notation 1.8.
@@
  • For ${\mathcal{O}}^{\otimes}={\mathrm{{\mathcal{F}}in}}_{*}$ we set $\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{D}}):=\mathrm{Alg}_{{\mathrm{{\mathcal{F}}in}}_{*}}({\mathcal{D}})$.
@@
  • If ${\mathcal{D}}^{\otimes}\to{\mathrm{{\mathcal{F}}in}}_{*}$ is a symmetric monoidal $\infty$-category, let ${\mathrm{coAlg}}_{{\mathbb{E}}_{\infty}}({\mathcal{D}}):=\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{D}}^{\mathrm{op}})^{\mathrm{op}}$.
2. $\infty$-operads and $\infty$-cooperads
@@ 2.1. The composition product
In this section we define the composition product following [3, Definition 3.3.].
@@
Let $\Sigma\simeq\coprod_{{\mathrm{n}}\geq 0}{\mathrm{B}}\Sigma_{\mathrm{n}}$ be the groupoid of finite sets and bijections.
@@
${\mathrm{sSeq}}({\mathcal{C}}):=\mathrm{Fun}(\Sigma,{\mathcal{C}})\simeq\prod_{{\mathrm{n}}\geq 0}\mathrm{Fun}({\mathrm{B}}\Sigma_{\mathrm{n}},{\mathcal{C}})$
@@
For every symmetric monoidal $\infty$-category ${\mathcal{C}}$ compatible with small colimits the $\infty$-category ${\mathrm{sSeq}}({\mathcal{C}})$ carries a symmetric monoidal structure compatible with small colimits given by Day-convolution [23, Proposition 4.8.1.10.], where we use the symmetric monoidal structure on $\Sigma$ taking disjoint union. For symmetric sequences ${\mathrm{X}},{\mathrm{Y}}$ in ${\mathcal{C}}$ we have
$({\mathrm{X}}\otimes{\mathrm{Y}})_{\mathrm{n}}:=\mathrm{colim}_{{\mathrm{i}}\coprod{\mathrm{j}}={\mathrm{n}}}{\mathrm{X}}_{\mathrm{i}}\otimes{\mathrm{Y}}_{\mathrm{j}}.$
@@
By [23, Proposition 4.8.1.17.] there is a symmetric monoidal equivalence ${\mathrm{sSeq}}({\mathcal{C}})\simeq{\mathcal{C}}\otimes{\mathrm{sSeq}}({\mathcal{S}})$ and $\iota$ is induced by the similarly defined embedding $\iota:{\mathcal{S}}\to{\mathrm{sSeq}}({\mathcal{S}}).$
@@
Observe that $\Sigma$ is the free symmetric monoidal $\infty$-category generated by a point. Thus ${\mathrm{sSeq}}({\mathcal{S}})\simeq\mathrm{Fun}(\Sigma^{\mathrm{op}},{\mathcal{S}})$ endowed with Day convolution is the free symmetric monoidal $\infty$-category compatible with small colimits generated by a point [23, Proposition 4.8.1.10.]. Tensoring with ${\mathcal{C}}$ we find that ${\mathcal{C}}\otimes{\mathrm{sSeq}}({\mathcal{S}})\simeq{\mathrm{sSeq}}({\mathcal{C}})$ is the free symmetric monoidal $\infty$-category compatible with small colimits under ${\mathcal{C}}$ generated by a point. So the composition
@@
By definition ${\mathrm{triv}}$ becomes the tensor unit of the composition product. For every ${\mathrm{X}}\in{\mathrm{sSeq}}({\mathcal{C}})$ there is a canonical equivalence ${\mathrm{X}}\simeq\coprod_{{\mathrm{k}}\geq 0}\iota({\mathrm{X}}_{\mathrm{k}})\otimes_{\Sigma_{\mathrm{k}}}{\mathrm{triv}}^{\otimes{\mathrm{k}}}.$
@@
$\Psi^{-1}({\mathrm{Y}})(\coprod_{{\mathrm{k}}\geq 0}\iota({\mathrm{X}}_{\mathrm{k}})\otimes_{\Sigma_{\mathrm{k}}}{\mathrm{triv}}^{\otimes{\mathrm{k}}})\simeq\coprod_{{\mathrm{k}}\geq 0}\iota({\mathrm{X}}_{\mathrm{k}})\otimes_{\Sigma_{\mathrm{k}}}{\mathrm{Y}}^{\otimes{\mathrm{k}}}.$
Thus for every ${\mathrm{n}}\in\Sigma$ there is an equivalence
$({\mathrm{X}}\circ{\mathrm{Y}})_{\mathrm{n}}\simeq\underset{{\mathrm{k}}\geq 0}{\coprod}(\underset{{{\mathrm{n}}_{1}\coprod...\coprod{\mathrm{n}}_{\mathrm{k}}={\mathrm{n}}}}{\mathrm{colim}}{\mathrm{X}}_{\mathrm{k}}\otimes(\bigotimes_{1\leq{\mathrm{j}}\leq{\mathrm{k}}}{\mathrm{Y}}_{{\mathrm{n}}_{\mathrm{j}}}))_{\Sigma_{\mathrm{k}}}.$
@@
$({\mathrm{X}}\circ{\mathrm{Y}})_{\mathrm{n}}\simeq\underset{(\mathsf{f}:{\mathrm{n}}\to{\mathrm{k}})\in({\mathrm{{\mathcal{F}}in}}_{{\mathrm{n}}/})^{\cong}}{\mathrm{colim}}{\mathrm{X}}_{\mathrm{k}}\otimes\bigotimes_{{\mathrm{j}}\in{\mathrm{k}}}{\mathrm{Y}}_{\mathsf{f}^{-1}({\mathrm{j}})}.$
The latter expression generalizes the following way: for $\ell\geq 2$ let ${\mathrm{{\mathcal{F}}in}}^{\ell}_{\mathrm{n}}$ be the groupoid whose objects are sequences of maps of finite sets ${\mathrm{n}}\xrightarrow{\mathsf{f}_{\ell-1}}{\mathrm{n}}_{\ell-1}\xrightarrow{\mathsf{f}_{\ell-2}}...\xrightarrow{\mathsf{f}_{1}}{\mathrm{n}}_{1}$ of length $\ell-1$ and the evident isomorphisms. For any ${\mathcal{O}}_{1},...,{\mathcal{O}}_{\ell}\in{\mathrm{sSeq}}({\mathcal{C}})$ there is an equivalence
$({\mathcal{O}}_{1}\circ...\circ{\mathcal{O}}_{\ell})_{\mathrm{n}}\simeq\mathrm{colim}_{\mathsf{f}\in{\mathrm{{\mathcal{F}}in}}^{\ell}_{\mathrm{n}}}(({\mathcal{O}}_{1})_{{\mathrm{n}}_{1}}\otimes\bigotimes_{{\mathrm{i}}\in{\mathrm{n}}_{1}}({\mathcal{O}}_{2})_{\mathsf{f}_{1}^{-1}({\mathrm{i}})}\otimes...\otimes\bigotimes_{{\mathrm{i}}\in{\mathrm{n}}_{\ell-1}}({\mathcal{O}}_{\ell})_{\mathsf{f}_{\ell-1}^{-1}({\mathrm{i}})})$
(see [6, Definition 2.12.]).
@@
Let $\Sigma_{\geq 1}:=\coprod_{{\mathrm{n}}\geq 1}{\mathrm{B}}\Sigma_{\mathrm{n}}$ and for every $\infty$-category ${\mathcal{C}}$ let
${\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}:=\mathrm{Fun}(\Sigma_{\geq 1},{\mathcal{C}})\simeq\prod_{{\mathrm{n}}\geq 1}\mathrm{Fun}({\mathrm{B}}\Sigma_{\mathrm{n}},{\mathcal{C}}).$
@@
${\mathrm{X}}\circ\iota({\mathrm{Y}})\simeq\coprod_{{\mathrm{k}}\geq 0}\iota({\mathrm{X}}_{\mathrm{k}})\otimes_{\Sigma_{\mathrm{k}}}\iota({\mathrm{Y}})^{\otimes{\mathrm{k}}}\simeq\iota(\coprod_{{\mathrm{k}}\geq 0}{\mathrm{X}}_{\mathrm{k}}\otimes_{\Sigma_{\mathrm{k}}}{\mathrm{Y}}^{\otimes{\mathrm{k}}})$
@@
Since the functor ${\mathcal{D}}\otimes_{\mathcal{C}}(-):\mathrm{Alg}_{{\mathbb{E}}_{\infty}}(\mathsf{Cat}_{\infty}^{{\mathrm{c}}{\mathrm{c}}})_{{\mathcal{C}}/}\to\mathrm{Alg}_{{\mathbb{E}}_{\infty}}(\mathsf{Cat}_{\infty}^{{\mathrm{c}}{\mathrm{c}}})_{{\mathcal{D}}/}$ is lax $\mathsf{Cat}_{\infty}^{{\mathrm{c}}{\mathrm{c}}}$-linear, it yields a monoidal functor on composition products:
@@
We will often consider the composition product in case that ${\mathcal{C}}$ has small colimits but the tensor product of ${\mathcal{C}}$ does not preserve small colimits component-wise. In this case the composition product does not define a monoidal $\infty$-category but a weaker structure based on the notion of non-symmetric $\infty$-operad [23, Definition 4.1.3.2.]:
@@
Let ${\mathcal{M}}^{\otimes}\to{\mathrm{BM}}^{\otimes}$ be a ${\mathrm{BM}}$-operad that exhibits an $\infty$-category ${\mathcal{D}}$ as weakly bitensored over lax monoidal $\infty$-categories ${\mathcal{C}},{\mathcal{E}}$. We say that ${\mathcal{M}}^{\otimes}\to{\mathrm{BM}}^{\otimes}$ exhibits ${\mathcal{D}}$ as lax bitensored over ${\mathcal{C}},{\mathcal{E}}$ if the functor ${\mathcal{M}}^{\otimes}\to{\mathrm{BM}}^{\otimes}$ is a locally cocartesian fibration and the restricted weak biaction of the monoidal subcategories of ${\mathcal{C}},{\mathcal{E}}$ spanned by the tensor unit is a bitensored $\infty$-category.
@@
Let ${\mathcal{V}}^{\otimes}\to{\mathrm{Ass}}$ be a non-symmetric $\infty$-operad and locally cocartesian fibration. Then for every ${\mathrm{V}}_{1},...,{\mathrm{V}}_{\mathrm{n}}$ there is a tensor product ${\mathrm{V}}_{1}\otimes...\otimes{\mathrm{V}}_{\mathrm{n}}\in{\mathcal{V}}$.
Moreover for every ${\mathrm{k}}\geq 0$, ${\mathrm{n}}_{1},...,{\mathrm{n}}_{\mathrm{k}}\geq 0$ and ${\mathrm{V}}^{\mathrm{i}}_{1},...,{\mathrm{V}}^{\mathrm{i}}_{{\mathrm{n}}_{\mathrm{i}}}\in{\mathcal{V}}$ for $1\leq{\mathrm{i}}\leq{\mathrm{k}}$ there is a canonical morphism in ${\mathcal{V}}:$
$\theta:{\mathrm{V}}^{1}_{1}\otimes...\otimes{\mathrm{V}}^{1}_{{\mathrm{n}}_{1}}\otimes...\otimes{\mathrm{V}}^{\mathrm{k}}_{1}\otimes...\otimes{\mathrm{V}}^{\mathrm{k}}_{{\mathrm{n}}_{\mathrm{k}}}\to({\mathrm{V}}^{1}_{1}\otimes...\otimes{\mathrm{V}}^{1}_{{\mathrm{n}}_{1}})\otimes...\otimes({\mathrm{V}}^{\mathrm{k}}_{1}\otimes...\otimes{\mathrm{V}}^{\mathrm{k}}_{{\mathrm{n}}_{\mathrm{k}}}).$
The functor ${\mathcal{V}}^{\otimes}\to{\mathrm{Ass}}$ is a monoidal $\infty$-category if and only if $\theta$ is an equivalence for all choices of inputs. The functor ${\mathcal{V}}^{\otimes}\to{\mathrm{Ass}}$ exhibits ${\mathcal{V}}$ as a lax monoidal $\infty$-category if and only if $\theta$ is an equivalence for ${\mathrm{V}}^{\mathrm{i}}_{1}=...={\mathrm{V}}^{\mathrm{i}}_{{\mathrm{n}}_{\mathrm{i}}}={\mathbb{1}}$ for $1\leq{\mathrm{i}}\leq{\mathrm{k}}$. In particular, for ${\mathrm{n}}_{1}=...={\mathrm{n}}_{\mathrm{k}}=0$ we find that the canonical morphism ${\mathbb{1}}\to{\mathbb{1}}^{\otimes{\mathrm{k}}}$ in ${\mathcal{V}}$ is an equivalence.
@@
    The tensor unit is ${\mathrm{triv}}$ and for every $\ell\geq 2$, any ${\mathcal{O}}_{1},...,{\mathcal{O}}_{\ell}\in{\mathrm{sSeq}}({\mathcal{C}})$ and ${\mathrm{n}}\in\Sigma$ there is an equivalence
    $({\mathcal{O}}_{1}\circ...\circ{\mathcal{O}}_{\ell})_{\mathrm{n}}\simeq\mathrm{colim}_{\mathsf{f}\in{\mathrm{{\mathcal{F}}in}}^{\ell}_{\mathrm{n}}}(({\mathcal{O}}_{1})_{{\mathrm{n}}_{1}}\otimes\bigotimes_{{\mathrm{i}}\in{\mathrm{n}}_{1}}({\mathcal{O}}_{2})_{\mathsf{f}_{1}^{-1}({\mathrm{i}})}\otimes...\otimes\bigotimes_{{\mathrm{i}}\in{\mathrm{n}}_{\ell-1}}({\mathcal{O}}_{\ell})_{\mathsf{f}_{\ell-1}^{-1}({\mathrm{i}})}).$

### v2

@@ 1. Introduction
(Theorem 4.40) Let $n$ be a natural and $X$ a $v_{n}$-periodic homotopy type. The unit
$\Phi(X)[-1]\to{\mathrm{Prim}}{\mathcal{U}}(\Phi(X)[-1])$
identifies with the Goodwillie completion $\Phi\to\lim_{n\geq 0}P_{n}(\Phi)$ evaluated at $\Omega(X)$.
In view of Theorem 1.3 the question arises which $v_{n}$-periodic homotopy types have a $\Phi$-good loop space. This is in general a hard question. However it is expected that any $v_{n}$-periodic homotopy type has a $\Phi$-good loop space for any prime and chromatic height $n$ (see [4] last paragraph). See [34] for examples.
We deduce Theorem 1.1 from the following theorem:
Theorem 1.4.
(Theorem 3.13) Let $\mathcal{C}$} be a stable presentably symmetric monoidal $\infty$-category, $\mathcal{O}$} a non-unital $\infty$-operad in $\mathcal{C}$} such that ${\mathcal{O}}_{1}={\mathbb{1}}$ and $\mathrm{X}$} an $\mathcal{O}$}-algebra in $\mathcal{C}$}. There is a commutative square:
[…]
To prove Theorem 1.4 we analyze Koszul duality and the interaction between the composition product and its dual version, the cocomposition product. Although Theorem 1.4 requires to assume that all norm maps of objects with an action of a symmetric group are invertible, a condition satisfied in all our examples of interest, we recently learned that this condition is not logically necessary but only necessary to apply our techniques: Heuts [31, Theorem 2.1.] recently proved a generalization of Theorem 1.4 without assuming this condition but using different techniques.
@@ 1.1. Acknowledgements
We thank Lukas Brantner, Gijs Heuts, Ben Knudsen, Markus Spitzweck and Manfred Stelzer for helpful discussions. We are especially grateful to Gijs Heuts for communicating to us the proof of Lemma 4.26. This work are main parts of my thesis written during my PhD 2014-2018 at the university of Osnabrück. I am grateful for this opportunity.
@@ 1.2. Notation and Terminologie
  • $\mathrm{Set}$ for the category of small sets.
  • $\Delta$ for the category of finite, non-empty, partially ordered sets and order preserving maps, whose objects we denote by ${n}]=\{0<...<{n}\}$ for ${n}\geq 0$.
  • $\mathrm{Fin}$} for (a skeleton of) the category of finite sets, whose objects we denote by ${n}=\{1,...,{n}\}$ for ${n}\geq 0$.
  • $\mathrm{Fin}$$_{*}$ for (a skeleton of) the category of finite pointed sets, whose objects we denote by $\langle{n}\rangle=\{*,1,...,{n}\}$ for ${n}\geq 0$.
@@
  • $\mathrm{Cat}_{\infty}$ for the $\infty$-category of small $\infty$-categories.
  • $\mathrm{Cat}_{\infty}^{{c}{c}}$ for the $\infty$-category of large $\infty$-categories with small colimits and small colimits preserving functors.
We often indicate $\infty$-categories of large objects by $\widehat{(-)}$, for example we write $\widehat{{\mathcal{S}}},\widehat{\mathrm{Cat}}_{\infty}$ for the $\infty$-categories of large spaces, $\infty$-categories.
@@
We heavily use symmetric $\infty$-operads [39, Definition 2.1.1.10.] and symmetric monoidal $\infty$-categories [39, Definition 2.0.0.7.]. For every $\infty$-operad ${\mathcal{O}}^{\otimes}\to{\mathrm{Fin}}_{*}$ we write $\mathcal{O}$} for the fiber over $\langle 1\rangle\in{\mathrm{Fin}}_{*}$ and call $\mathcal{O}$} the underlying $\infty$-category. More generally, we use $\mathcal{O}$}-operads and $\mathcal{O}$}-monoidal $\infty$-categories [39, Definition 2.1.2.13.], which for ${\mathcal{O}}^{\otimes}={\mathrm{Fin}}_{*}$ give symmetric $\infty$-operads and symmetric monoidal $\infty$-categories, for ${\mathcal{O}}^{\otimes}={\mathrm{Ass}}$ give non-symmetric $\infty$-operads and monoidal $\infty$-categories and for ${\mathcal{O}}^{\otimes}={\mathrm{LM}},{\mathrm{RM}},{\mathrm{BM}}$ give ${\mathrm{LM}},{\mathrm{RM}},{\mathrm{BM}}$-operads and left, right, bitensored $\infty$-categories.
Notation 1.5.
@@
  • For ${\mathcal{O}}^{\otimes}={\mathrm{Fin}}_{*}$ we set $\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{D}}):=\mathrm{Alg}_{{\mathrm{Fin}}_{*}}({\mathcal{D}})$.
@@
  • If ${\mathcal{D}}^{\otimes}\to{\mathrm{Fin}}_{*}$ is a symmetric monoidal $\infty$-category, let $\mathrm{co}\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{D}}):=\mathrm{Alg}_{{\mathbb{E}}_{\infty}}({\mathcal{D}}^{\mathrm{op}})^{\mathrm{op}}$.
2. $\infty$-operads
@@ 2.1. The composition product
In this section we define the composition product following [8, Definition 3.3.].
@@
Let $\Sigma\simeq\coprod_{{n}\geq 0}{\mathrm{B}}\Sigma_{n}$ be the groupoid of finite sets and bijections.
@@
${\mathrm{sSeq}}({\mathcal{C}}):=\mathrm{Fun}(\Sigma,{\mathcal{C}})\simeq\prod_{{n}\geq 0}\mathrm{Fun}({\mathrm{B}}\Sigma_{n},{\mathcal{C}})$
@@
For every symmetric monoidal $\infty$-category $\mathcal{C}$} compatible with small colimits the $\infty$-category ${\mathrm{sSeq}}({\mathcal{C}})$ carries a symmetric monoidal structure compatible with small colimits given by Day-convolution [39, Proposition 4.8.1.10.], where we use the symmetric monoidal structure on $\Sigma$ taking disjoint union. For symmetric sequences ${\mathrm{X}},{\mathrm{Y}}$ in $\mathcal{C}$} we have
$({\mathrm{X}}\otimes{\mathrm{Y}})_{n}:=\mathrm{colim}_{{i}\coprod{j}={n}}{\mathrm{X}}_{i}\otimes{\mathrm{Y}}_{j}.$
@@
By [39, Proposition 4.8.1.17.] there is a symmetric monoidal equivalence ${\mathrm{sSeq}}({\mathcal{C}})\simeq{\mathcal{C}}\otimes{\mathrm{sSeq}}({\mathcal{S}})$ and $\iota$ is induced by the similarly defined embedding $\iota:{\mathcal{S}}\to{\mathrm{sSeq}}({\mathcal{S}}).$
@@
Observe that $\Sigma$ is the free symmetric monoidal $\infty$-category generated by a point. Thus ${\mathrm{sSeq}}({\mathcal{S}})\simeq\mathrm{Fun}(\Sigma^{\mathrm{op}},{\mathcal{S}})$ endowed with Day convolution is the free symmetric monoidal $\infty$-category compatible with small colimits generated by a point [39, Proposition 4.8.1.10.]. Tensoring with $\mathcal{C}$} we find that ${\mathcal{C}}\otimes{\mathrm{sSeq}}({\mathcal{S}})\simeq{\mathrm{sSeq}}({\mathcal{C}})$ is the free symmetric monoidal $\infty$-category compatible with small colimits under $\mathcal{C}$} generated by a point. So the composition
@@
By definition $\mathrm{triv}$} becomes the tensor unit of the composition product. For every ${\mathrm{X}}\in{\mathrm{sSeq}}({\mathcal{C}})$ there is a canonical equivalence ${\mathrm{X}}\simeq\coprod_{{k}\geq 0}\iota({\mathrm{X}}_{k})\otimes_{\Sigma_{k}}{\mathrm{triv}}^{\otimes{k}}.$
@@
$\Psi^{-1}({\mathrm{Y}})(\coprod_{{k}\geq 0}\iota({\mathrm{X}}_{k})\otimes_{\Sigma_{k}}{\mathrm{triv}}^{\otimes{k}})\simeq\coprod_{{k}\geq 0}\iota({\mathrm{X}}_{k})\otimes_{\Sigma_{k}}{\mathrm{Y}}^{\otimes{k}}.$
Thus for every ${n}\in\Sigma$ there is an equivalence
$({\mathrm{X}}\circ{\mathrm{Y}})_{n}\simeq\underset{{k}\geq 0}{\coprod}(\underset{{{n}_{1}\coprod...\coprod{n}_{k}={n}}}{\mathrm{colim}}{\mathrm{X}}_{k}\otimes(\bigotimes_{1\leq{j}\leq{k}}{\mathrm{Y}}_{{n}_{j}}))_{\Sigma_{k}}.$
@@
$({\mathrm{X}}\circ{\mathrm{Y}})_{n}\simeq\underset{({\mathrm{f}}:{n}\to{k})\in({\mathrm{Fin}}_{{n}/})^{\cong}}{\mathrm{colim}}{\mathrm{X}}_{k}\otimes\bigotimes_{{j}\in{k}}{\mathrm{Y}}_{{\mathrm{f}}^{-1}({j})}.$
The latter expression generalizes the following way: for $\ell\geq 2$ let ${\mathrm{Fin}}^{\ell}_{n}$ be the groupoid whose objects are sequences of maps of finite sets ${n}\xrightarrow{{\mathrm{f}}_{\ell-1}}{n}_{\ell-1}\xrightarrow{{\mathrm{f}}_{\ell-2}}...\xrightarrow{{\mathrm{f}}_{1}}{n}_{1}$ of length $\ell-1$ and the evident isomorphisms. For any ${\mathcal{O}}_{1},...,{\mathcal{O}}_{\ell}\in{\mathrm{sSeq}}({\mathcal{C}})$ there is an equivalence
$({\mathcal{O}}_{1}\circ...\circ{\mathcal{O}}_{\ell})_{n}\simeq\mathrm{colim}_{{\mathrm{f}}\in{\mathrm{Fin}}^{\ell}_{n}}(({\mathcal{O}}_{1})_{{n}_{1}}\otimes\bigotimes_{{i}\in{n}_{1}}({\mathcal{O}}_{2})_{{\mathrm{f}}_{1}^{-1}({i})}\otimes...\otimes\bigotimes_{{i}\in{n}_{\ell-1}}({\mathcal{O}}_{\ell})_{{\mathrm{f}}_{\ell-1}^{-1}({i})})$
(see [12, Definition 2.12.]).
@@
Let $\Sigma_{\geq 1}:=\coprod_{{n}\geq 1}{\mathrm{B}}\Sigma_{n}$ and for every $\infty$-category $\mathcal{C}$} let
${\mathrm{sSeq}}({\mathcal{C}})_{\geq 1}:=\mathrm{Fun}(\Sigma_{\geq 1},{\mathcal{C}})\simeq\prod_{{n}\geq 1}\mathrm{Fun}({\mathrm{B}}\Sigma_{n},{\mathcal{C}}).$
@@
${\mathrm{X}}\circ\iota({\mathrm{Y}})\simeq\coprod_{{k}\geq 0}\iota({\mathrm{X}}_{k})\otimes_{\Sigma_{k}}\iota({\mathrm{Y}})^{\otimes{k}}\simeq\iota(\coprod_{{k}\geq 0}{\mathrm{X}}_{k}\otimes_{\Sigma_{k}}{\mathrm{Y}}^{\otimes{k}})$
@@
Since the functor ${\mathcal{D}}\otimes_{\mathcal{C}}(-):\mathrm{Alg}_{{\mathbb{E}}_{\infty}}(\mathrm{Cat}_{\infty}^{{c}{c}})_{{\mathcal{C}}/}\to\mathrm{Alg}_{{\mathbb{E}}_{\infty}}(\mathrm{Cat}_{\infty}^{{c}{c}})_{{\mathcal{D}}/}$ is lax $\mathrm{Cat}_{\infty}^{{c}{c}}$-linear, it yields a monoidal functor on composition products:
@@
We will often consider the composition product in case that $\mathcal{C}$} has small colimits but the tensor product of $\mathcal{C}$} does not preserve small colimits component-wise. In this case the composition product does not define a monoidal $\infty$-category but a weaker structure based on the notion of non-symmetric $\infty$-operad [39, Definition 4.1.3.2.]:
@@
Let ${\mathcal{M}}^{\otimes}\to{\mathrm{BM}}^{\otimes}$ be a $\mathrm{BM}$}-operad that exhibits an $\infty$-category $\mathcal{D}$} as weakly bitensored over lax monoidal $\infty$-categories $\mathcal{C}$$\mathcal{E}$}. We say that ${\mathcal{M}}^{\otimes}\to{\mathrm{BM}}^{\otimes}$ exhibits $\mathcal{D}$} as lax bitensored over $\mathcal{C}$$\mathcal{E}$} if the functor ${\mathcal{M}}^{\otimes}\to{\mathrm{BM}}^{\otimes}$ is a locally cocartesian fibration and the restricted weak biaction of the monoidal subcategories of $\mathcal{C}$$\mathcal{E}$} spanned by the tensor unit is a bitensored $\infty$-category.
@@
Let ${\mathcal{V}}^{\otimes}\to{\mathrm{Ass}}$ be a non-symmetric $\infty$-operad and locally cocartesian fibration. Then for every ${\mathrm{V}}_{1},...,{\mathrm{V}}_{n}$ there is a tensor product ${\mathrm{V}}_{1}\otimes...\otimes{\mathrm{V}}_{n}\in{\mathcal{V}}$.
Moreover for every ${k}\geq 0$, ${n}_{1},...,{n}_{k}\geq 0$ and ${\mathrm{V}}^{i}_{1},...,{\mathrm{V}}^{i}_{{n}_{i}}\in{\mathcal{V}}$ for $1\leq{i}\leq{k}$ there is a canonical morphism in $\mathcal{V}$}:
$\theta:{\mathrm{V}}^{1}_{1}\otimes...\otimes{\mathrm{V}}^{1}_{{n}_{1}}\otimes...\otimes{\mathrm{V}}^{k}_{1}\otimes...\otimes{\mathrm{V}}^{k}_{{n}_{k}}\to({\mathrm{V}}^{1}_{1}\otimes...\otimes{\mathrm{V}}^{1}_{{n}_{1}})\otimes...\otimes({\mathrm{V}}^{k}_{1}\otimes...\otimes{\mathrm{V}}^{k}_{{n}_{k}}).$
The functor ${\mathcal{V}}^{\otimes}\to{\mathrm{Ass}}$ is a monoidal $\infty$-category if and only if $\theta$ is an equivalence for all choices of inputs. The functor ${\mathcal{V}}^{\otimes}\to{\mathrm{Ass}}$ exhibits $\mathcal{V}$} as a lax monoidal $\infty$-category if and only if $\theta$ is an equivalence for ${\mathrm{V}}^{i}_{1}=...={\mathrm{V}}^{i}_{{n}_{i}}={\mathbb{1}}$ for $1\leq{i}\leq{k}$. In particular, for ${n}_{1}=...={n}_{k}=0$ we find that the canonical morphism ${\mathbb{1}}\to{\mathbb{1}}^{\otimes{k}}$ in $\mathcal{V}$} is an equivalence.
@@
    The tensor unit is $\mathrm{triv}$} and for every $\ell\geq 2$, any ${\mathcal{O}}_{1},...,{\mathcal{O}}_{\ell}\in{\mathrm{sSeq}}({\mathcal{C}})$ and ${n}\in\Sigma$ there is an equivalence
    $({\mathcal{O}}_{1}\circ...\circ{\mathcal{O}}_{\ell})_{n}\simeq\mathrm{colim}_{{\mathrm{f}}\in{\mathrm{Fin}}^{\ell}_{n}}(({\mathcal{O}}_{1})_{{n}_{1}}\otimes\bigotimes_{{i}\in{n}_{1}}({\mathcal{O}}_{2})_{{\mathrm{f}}_{1}^{-1}({i})}\otimes...\otimes\bigotimes_{{i}\in{n}_{\ell-1}}({\mathcal{O}}_{\ell})_{{\mathrm{f}}_{\ell-1}^{-1}({i})}).$